OpenStudy (anonymous):

bob bought 3 one third yards of fabric. if he use 2/3 of it how much does she have left??

OpenStudy (anonymous):

3 1/2

OpenStudy (anonymous):

1yard = 3foot, --> 3\(\dfrac{1}{3}\)yard =\(\dfrac{10}{3}yard \)= 10 foot now, \(\dfrac{2}{3}*10 = 20/3 foot,\) so the leftover is 10-20/3 = 10/3 foot (not yard)

OpenStudy (anonymous):

its 3 1/2 sorry